Emily rows 6 miles downstream in 1 hour and Ashley, rowing 1 mile per hour faster, completes the return trip in 2 hours.
1. Find the speed of the current (c) and each girl's rowing speed.
2. If Emily and Ashley were rowing separately, who would complete their trip first and by how long? Round to the nearest hundredth if necessary.

Answer by lwsshak3(11628)

Emily rows 6 miles downstream in 1 hour and Ashley, rowing 1 mile per hour faster, completes the return trip in 2 hours.
1. Find the speed of the current (c) and each girl's rowing speed.
2. If Emily and Ashley were rowing separately, who would complete their trip first and by how long? Round to the nearest hundredth if necessary.
***
let c=speed of current
let x=Emily's rowing speed
x+c=Emily's speed downstream
x+1=Ashley's rowing speed
x+1-c=Ashley's speed upstream
travel time*speed=distance
..
1*(x+c)=6
2*(x+1-c)=6
..
x+c=6
2x+2-2c=6
..
x+c=6
2x-2c=4
..
2x+2c=12
2x-2c=4
add
4x=16
x=4
c=6-x=2
speed of current=2 mph
Emily's rowing speed=4 mph
Ashley's rowing speed=5 mph
rowing separately downstream and upstream:
6/(4+2)+6/(4-2)=3 hrs (Emily's travel time)
6/(5+2)+6/(5-2)=2-6/7 hr (Ashley's travel time)
Ashley completes her trip in 1/7 hr less time than Emily